OpenStudy (anonymous):

What is the exact value of cos(2 x) if tan( x) = -3/4 and x is in quadrant IV?

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

how did you come up wit that?

OpenStudy (ankit042):

1+tan^2 x = sec^2 x have you seen this formula?

OpenStudy (anonymous):

no I havent

OpenStudy (ankit042):

K there are three trigonometric identities this is one of them. Can you solve it now?

OpenStudy (ankit042):

also I think -1/2 is not the correct answer. As it is in 4th quad cos is always +ve

OpenStudy (ankit042):

there is this direct formula which you can use cos 2x = (1 - tan^2 x) / (1 + tan^2 x)

OpenStudy (dumbcow):

 from the triangle in 4th quadrant you can get sin and cos \[\sin x = -\frac{3}{5}\] \[\cos x = \frac{4}{5}\] the double angle identity is: \[\cos 2x = \cos^2 x - \sin^2 x\]

OpenStudy (anonymous):

ok so cos is 3/2?

OpenStudy (anonymous):

wait no im lost

OpenStudy (ankit042):

cos is always between +1 and -1 lol

OpenStudy (ankit042):

cos2x = 16/25-9/25

OpenStudy (anonymous):

7/25= cos2x?